\begin{thebibliography}{99}                                                                                               %


\bibitem {BCHMP}Briand, P., Coquet, F., Hu, Y., M\'{e}min, J., Peng, S., 2000.
A converse comparison theorem for BSDEs and related properties of
g-expectation. Electron. Comm. Probab. 5, 101-117.

\bibitem {CCD}Chen, Z.J., Chen, T., Davison, M., 2005. Choquet expectation
and Peng's $g-$expectation. The Annals of Probability 33(3), 1179-1199.

\bibitem {CKW}Chen, Z.J., Kulperger, R., Wei, G., 2005. A comonotonic theorem
for BSDEs. Stochastic Processes and their Applications 115, 41-54.

\bibitem {CS}Chen, Z.J., Sulem, A., 2001. An integral representation theorem
of $g$-expectations. Research Report INRIA, No.4284.

\bibitem {Choq}Choquet, G., 1953. Theory of capacities. Ann. Inst. Fourier
(Grenoble) 5, 131-195.

\bibitem {CHMP}Coquet, F., Hu, Y., M\'{e}min, J., Peng, S., 2002. Filtration
consistent nonlinear expectations and related g-expectations. Probab. Theory
and Related Fields 123, 1-27.

\bibitem {De}Dellacherie, C., 1991. Quelques commentaires sur les
prolongements de capacit\'{e}s. In: Strasbourg, V.(Ed.). Seminaire de
probabilites. Springer, Berlin, 77-81.

\bibitem {Den}Denneberg, D., 1994. Non-additive Measure and Integral. Kluwer
Academic Publishers, Boston.

\bibitem {EPQ}El Karoui, N., Peng, S., Quenez, M.C., 1997. Backward stochastic
differential equations in finance. Math. Finance 7, 1-71.

\bibitem {HHC}He, K., Hu, M., Chen, Z.J., 2009. The relationship between risk
measures and Choquet expectations in the framework of $g$-expectations.
Statistics and Probability Letters 79, 508-512.

\bibitem {J1}Jiang, L., 2006. A note on $g$-expectation with
comonotonic additivity. Statistics and Probability Letters 76,
1895-1903.

\bibitem {J}Jiang, L., 2008. Convexity, translation invariance and
subadditivity for $g$-expectations and related risk measures. Annals of
Applied Probability 18(1), 245-258.

\bibitem {PP}Pardoux, E., Peng, S., 1990. Adapted solution of a backward
stochastic differential equation. Systems and Control Letters 14, 55-61.

\bibitem {P1}Peng, S., 1997. BSDE and Stochastic Optimizations. Topics in
Stochastic Analysis, Lecture Notes of 1995 Summer School in Math. Yan, J.,
Peng, S., Fang, S., Wu, L.M. Ch.2, (Chinese vers.), Science Press, Beijing.

\bibitem {P2}Peng, S., 1997. Backward SDE and related $g$-expectations.
Backward stochastic differential equations, in El N. Karoui and L. Mazliak,
eds. Pitman Res. Notes Math. Ser. Longman Harlow, vol. 364, 141-159.

\bibitem {S1}Schmeidler, D., 1986. Integral representation without additivity.
Proceedings of the American Mathematical Society 97, 255-261.
\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphicx}%
\providecommand{\U}[1]{\protect \rule{.1in}{.1in}}
\newtheorem{theorem}{Theorem}

\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{definition}[theorem]{Definition}

\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{remark}[theorem]{Remark}

\newenvironment{proof}[1][Proof]{\noindent \textbf{#1.} }{\  $\Box$}

\title{Choquet expectations and $g$-expectations with multi-dimensional
Brownian motion}

\author{Mingshang Hu\footnote{Email: humingshang\symbol{64}gmail.com}\\
        \footnotesize{
        School of Mathematics, Shandong University, Jinan, \small{250100}, China
 } }

\begin{document}
\maketitle

\noindent \textbf{Abstract. }We prove that a $g$-expectation is a
Choquet expectation if and only if $g$ is independent of $y$ and is
linear in $z$, i.e., classical linear expectation, without the
assumptions that the deterministic generator $g$ is continuous in
$t$ and the dimension of the Brownian motion is one.

\  \  \

\noindent \textbf{Keywords: }BSDE, $g$-expectation, conditional
$g$-expectation, capacity, Choquet expectation, comonotonic
additivity.

\section{Introduction}

Choquet \cite{Choq} introduced the notion of Choquet expectations
via capacities in 1953. Peng \cite{P2} introduced the notions of
$g$-expectations and conditional $g$-expectations via a class of
backward stochastic differential equations (BSDEs for short) in
1997. These two types of nonlinear mathematical expectations have
their own characteristics. For example, Choquet expectations are
comonotonic additivity, $g$-expectations and conditional
$g$-expectations are consistent. In Chen et al. \cite{CCD}, the
authors studied an interesting problem:

If a $g$-expectation is a Choquet expectation, can we find the form
of the generator $g$?

Under the assumptions that the deterministic generator $g$ is
continuous in $t$ and the dimension of the Brownian motion is one,
Chen et al. \cite{CCD} proved that a $g$-expectation is a Choquet
expectation if and only if $g$ is independent of $y$ and is linear
in $z$. For the case that the dimension of the Brownian motion is
greater than one, the main difficulty is to find the form of the
generator $g$. Unfortunately, this problem is not a simple extension
of the one dimensional case. Take a $2$-dimensional Brownian motion
$W_{t}=(W_t^1,W_t^2)$ for example, $W_t^1$ and $W_t^2$ are not
comonotonic. This prevents us from using the method in Chen et al.
\cite{CCD} directly. To overcome this defect, we consider
comonotonic indicator functions and use a property of BSDE.
Furthermore, our method does not need the continuous assumption on
$g$.

This paper is organized as follows: In Section 2, we recall some facts about
$g$-expectations and Choquet expectations. In Section 3, we state and prove
our main result.

\section{Preliminaries}

Fix $T>0$, let $(W_{t})_{0\leq t\leq T}$ be a $d$-dimensional standard
Brownian motion defined on a completed probability space $(\Omega
,\mathcal{F},P)$ and $(\mathcal{F}_{t})_{0\leq t\leq T}$ be the natural
filtration generated by this Brownian motion. For $x=(x_{1},\ldots,x_{d})$,
$y=(y_{1},\ldots,y_{d})\in \mathbb{R}^{d}$, $|x|:=\sqrt{\sum_{i=1}^{d}%
|x_{i}|^{2}}$, $x\cdot y:=\sum_{i=1}^{d}x_{i}y_{i}$. We denote by
$L^{2}(\mathcal{F}_{t})$ the set of all square integrable $\mathcal{F}_{t}%
$-measurable random variables and $L^{2}(0,T;\mathbb{R}^{n})$ the space of all
$\mathcal{F}_{t}$-adapted, $\mathbb{R}^{n}$-valued processes $(v_{t}%
)_{t\in \lbrack0,T]}$ with $E\int_{0}^{T}|v_{t}|^{2}dt<\infty$.

Let us consider a deterministic function $g:[0,T]\times \mathbb{R\times R}%
^{d}\rightarrow \mathbb{R}$, which will be in the following the generator of
the BSDE. For the function $g$, we will use the following assumptions:

\begin{description}
\item[(H1)] For each $(y,z)\in \mathbb{R\times R}^{d}$, $t\rightarrow g(t,y,z)$
is measurable.

\item[(H1$^{\prime}$)] For each $(y,z)\in \mathbb{R\times R}^{d}$,
$t\rightarrow g(t,y,z)$ is continuous.

\item[(H2)] There exists a constant $K\geq0$ such that
\[
|g(t,y,z)-g(t,y^{\prime},z^{\prime})|\leq K(|y-y^{\prime}|+|z-z^{\prime
}|),\  t\in \lbrack0,T],y,y^{\prime}\in \mathbb{R},z,z^{\prime}%
\in \mathbb{ R}^{d}.
\]

\item[(H3)] $g(t,y,0)\equiv0$ for each $(t,y)\in \lbrack0,T]\times \mathbb{R}.$
\end{description}

Let $g$ satisfy (H1)-(H3). Then for each $\xi \in L^{2}(\mathcal{F}_{T})$, the
BSDE%
\begin{equation}\label{e1}
y_{t}=\xi+\int_{t}^{T}g(s,y_{s},z_{s})ds-\int_{t}^{T}z_{s}\cdot dW_{s},\ 0\leq
t\leq T,
\end{equation}
has a unique solution $(y_{t,}z_{t})_{t\in \lbrack0,T]}\in L^{2}(0,T;\mathbb{R}%
)\times L^{2}(0,T;\mathbb{R}^{d})$ (see Pardoux and Peng \cite{PP}),
which depends on the generator $g$ and terminal value $\xi.$

The following standard estimate for BSDEs can be found in \cite{EPQ,
P1, BCHMP}.

\begin{lemma}\label{le1}
Suppose $g$ satisfies (H1)-(H3). For each $\xi^{1},\xi^{2}\in L^{2}%
(\mathcal{F}_{T})$, let $(y_{t,}^{i}z_{t}^{i})_{t\in \lbrack0,T]}$
be the solution of BSDE (\ref{e1}) corresponding to the generator
$g$ and terminal value
$\xi^{i}$ with $i=1,2.$ Then there exists a constant $C>0$ such that%
\[
E[\sup_{t\leq s\leq T}|y_{s}^{1}-y_{s}^{2}|^{2}|\mathcal{F}_{t}]+E[\int
_{t}^{T}|z_{s}^{1}-z_{s}^{2}|^{2}ds|\mathcal{F}_{t}]\leq CE[|\xi^{1}-\xi
^{2}|^{2}|\mathcal{F}_{t}].
\]

\end{lemma}

Using the solution of BSDE (\ref{e1}), Peng \cite{P2} proposed the
following notions:

\begin{definition}\label{de2}
Suppose $g$ satisfies (H1)-(H3). For each $\xi \in L^{2}(\mathcal{F}_{T}),$ let
$(y_{t,}z_{t})_{t\in \lbrack0,T]}$ be the solution of BSDE (\ref{e1}), define%
\[
\mathcal{E}_{g}[\xi]:=y_{0};\qquad \mathcal{E}_{g}[\xi|\mathcal{F}_{t}%
]:=y_{t}\quad \mbox{for each}\ t\in \lbrack0,T].
\]
$\mathcal{E}_{g}[\xi]$ is called the $g$-expectation of $\xi$ and
$\mathcal{E}_{g}[\xi|\mathcal{F}_{t}]$ is called the conditional
$g$-expectation of $\xi$ with respect to $\mathcal{F}_{t}$.
\end{definition}

We now recall the notions of capacity and Choquet expectation. A
capacity is a set function $V:\mathcal{F}_{T}\mapsto \lbrack0,1]$
satisfying: (i) $V(\emptyset)=0,$ $V(\Omega)=1;$ (ii) $V(A)\leq
V(B)$ for each $A\subset B.$ The corresponding Choquet expectation
(see \cite{Choq}) is defined as
follows:%
\[
\mathcal{C}[\xi]:=\int_{-\infty}^{0}[V(\xi \geq t)-1]dt+\int_{0}^{\infty}%
V(\xi \geq t)dt\quad \mbox{for each}\  \xi \in L^{2}(\mathcal{F}_{T}).
\]
Two random variables $\xi$ and $\eta$ are called comonotonic if%
\[
\lbrack \xi(\omega)-\xi(\omega^{\prime})][\eta(\omega)-\eta(\omega^{\prime
})]\geq0\quad \mbox{for each}\  \omega,\omega^{\prime}\in \Omega.
\]
Now, we list some properties of Choquet expectations (see
\cite{Choq, S1, De, Den}).

\begin{description}
\item[(1)] Monotonicity: If $\xi \geq \eta,$ then $\mathcal{C}[\xi
]\geq \mathcal{C}[\eta].$

\item[(2)] Positive homogeneity: If $\lambda \geq0,$ then $\mathcal{C}%
[\lambda \xi]=\lambda \mathcal{C}[\xi].$

\item[(3)] Translation invariance: If $c\in \mathbb{R},$ then $\mathcal{C}%
[\xi+c]=\mathcal{C}[\xi]+c.$

\item[(4)] Comonotonic additivity: If $\xi$ and $\eta$ are comonotonic, then
$\mathcal{C}[\xi+\eta]=\mathcal{C}[\xi]+\mathcal{C}[\eta].$
\end{description}

Let $g$ satisfy (H1)-(H3), define%
\[
P_{g}(A):=\mathcal{E}_{g}[I_{A}]\quad \mbox{for each}\ A\in \mathcal{F}_{T}.
\]
$P_{g}(A)$ is called the $g$-probability of $A.$ Obviously, $P_{g}(\cdot)$ is
a capacity. The corresponding Choquet expectation is denoted by $\mathcal{C}%
_{g}$. It is easy to check that $\mathcal{C}_{g}[I_{A}]=\mathcal{E}_{g}%
[I_{A}]$ for each $A\in \mathcal{F}_{T}$. Furthermore, $\mathcal{C}_{g}%
[\xi]<\infty$ for each $\xi \in L^{2}(\mathcal{F}_{T})$ (see
\cite{HHC}).

The following result can be found in \cite{CCD}.

\begin{lemma}\label{le3}
Suppose that $d=1$ and $g$ satisfies (H1$^{\prime}$), (H2) and (H3). Then
$\mathcal{E}_{g}[\xi]=\mathcal{C}_{g}[\xi]$ for each $\xi \in L^{2}%
(\mathcal{F}_{T})$ if and only if $g$ is independent of $y$ and is linear in
$z$, i.e., $g(t,z)=g(t,1)z$.
\end{lemma}

\section{Main result}

Let $\{e_{1},e_{2},\ldots,e_{d}\}$ denote the standard basis of $\mathbb{R}%
^{d}$. Now we give the main result.

\begin{theorem}\label{th4}
Suppose $g$ satisfies (H1)-(H3). Then $\mathcal{E}_{g}[\xi]=\mathcal{C}%
_{g}[\xi]$ for each $\xi \in L^{2}(\mathcal{F}_{T})$ if and only if $g$ is
independent of $y$ and is linear in $z$, i.e., $g(t,z)=\sum_{i=1}^{d}%
g(t,e_{i})z_{i}$ for almost every $t\in \lbrack0,T]$, where $z_{i}$ is the
$i$-th component of $z$.
\end{theorem}

For proving this theorem, we need the following lemmas. The first
lemma is a direct consequence of Jiang \cite{J} (see also
\cite{BCHMP, CCD, J1}).

\begin{lemma}\label{le5}
Suppose $g$ satisfies (H1)-(H3). If $\mathcal{E}_{g}[\xi
]=\mathcal{C}_{g}[\xi]$ for each $\xi \in L^{2}(\mathcal{F}_{T})$,
then $g$ is independent of $y$ and is positively homogeneous in $z.$
\end{lemma}

\begin{proof}
Since $\mathcal{E}_{g}=\mathcal{C}_{g},$ we have%
\[
\mathcal{E}_{g}[\xi+c]=\mathcal{E}_{g}[\xi]+c\quad \mbox{for
each}\ c\in \mathbb{R}%
;\  \mathcal{E}_{g}[\lambda \xi]=\lambda \mathcal{E}_{g}[\xi]\quad
\mbox{for each}\  \lambda \geq0.
\]
From this, we obtain the result (see Theorems 3.1 and 3.4 in Jiang
\cite{J}). The proof is complete.
\end{proof}

The next lemma is a property of BSDE (see \cite{P1}).

\begin{lemma}\label{le6}
Suppose $g$ satisfies (H1)-(H3). Let $\xi$ be a $k_{1}$-dimensional
$\mathcal{F}_{t_{0}}$-measurable random vector and $\eta$ be a $k_{2}%
$-dimensional $\mathcal{F}_{T}$-measurable random vector, where $t_{0}%
\in \lbrack0,T)$ and $k_{1},k_{2}\in \mathbb{N}$. Then for each $f\in
C_{b}(\mathbb{R}^{k_{1}}\times \mathbb{R}^{k_{2}})$, we have
\[
\mathcal{E}_{g}[f(\xi,\eta)|\mathcal{F}_{t}]=\mathcal{E}_{g}[f(x,\eta
)|\mathcal{F}_{t}]|_{x=\xi},\ t\in \lbrack t_{0},T].
\]

\end{lemma}

\begin{proof}
We outline the proof for the convenience of the reader. The proof is divided
into two steps.

Step 1: Let $\xi$ be simple random vector, i.e., $\xi=\sum_{i=1}^{N}%
x_{i}I_{A_{i}}$, where $\{x_{i}\}_{i=1}^{N}\subset
\mathbb{R}^{k_{1}}$ and $\{A_{i}\}_{i=1}^{N}$ is an
$\mathcal{F}_{t_{0}}$-partition of $\Omega$. Let
$(y_{t}^{i},z_{t}^{i})_{t\in \lbrack0,T]}$ denote the solution of
BSDE (\ref{e1}) corresponding to the generator $g$ and terminal
value $f(x_{i},\eta)$ with
$i=1,\ldots,N$. Then it is easy to verify that $(\sum_{i=1}^{N}y_{t}%
^{i}I_{A_{i}},\sum_{i=1}^{N}z_{t}^{i}I_{A_{i}})_{t\in \lbrack
t_{0},T]}$ is the solution of BSDE (\ref{e1}) on $[t_{0},T]$
corresponding to the generator $g$ and terminal value
$\sum_{i=1}^{N}f(x_{i},\eta)I_{A_{i}}$. Noting that
$f(\sum_{i=1}^{N}x_{i}I_{A_{i}},\eta)=\sum_{i=1}^{N}f(x_{i},\eta)I_{A_{i}}$,
then for $t\in \lbrack t_{0},T]$, we have%
\[
\mathcal{E}_{g}[f(\xi,\eta)|\mathcal{F}_{t}]=\sum_{i=1}^{N}\mathcal{E}%
_{g}[f(x_{i},\eta)|\mathcal{F}_{t}]I_{A_{i}}=\mathcal{E}_{g}[f(x,\eta
)|\mathcal{F}_{t}]|_{x=\xi}.
\]

Step 2: For general $\xi$, we can choose some simple random vectors $\xi
_{n}\rightarrow \xi$. Since $f\in C_{b}(\mathbb{R}^{k_{1}}\times \mathbb{R}%
^{k_{2}})$, by Lemma \ref{le1}, we get for $t\in \lbrack t_{0},T]$,%
\[
P-a.s.,\  \mathcal{E}_{g}[f(\xi_{n},\eta)|\mathcal{F}_{t}]\rightarrow
\mathcal{E}_{g}[f(\xi,\eta)|\mathcal{F}_{t}],\mathcal{E}_{g}[f(x,\eta
)|\mathcal{F}_{t}]|_{x=\xi_{n}}\rightarrow \mathcal{E}_{g}[f(x,\eta
)|\mathcal{F}_{t}]|_{x=\xi}.
\]
Thus $\mathcal{E}_{g}[f(\xi,\eta)|\mathcal{F}_{t}]=\mathcal{E}_{g}%
[f(x,\eta)|\mathcal{F}_{t}]|_{x=\xi}$. The proof is complete.
\end{proof}

\begin{remark}\label{re7}
Let $f_{n}\in C_{b}(\mathbb{R}^{k_{1}}\times \mathbb{R}^{k_{2}})$ be
uniformly bounded such that $f_{n}\rightarrow f$. Then by Lemma
\ref{le1}, we can easily prove that Lemma \ref{le6} still holds for
$f$.
\end{remark}

The following lemma plays an important role in proving the main theorem with
$d=1.$

\begin{lemma}\label{le8}
Suppose that $d=1$ and $g$ satisfies (H1)-(H3). If $\mathcal{E}_{g}%
[\xi]=\mathcal{C}_{g}[\xi]$ for each $\xi \in L^{2}(\mathcal{F}_{T})$, then for
each $t\in \lbrack0,T],$ $n\in \mathbb{N}$, we have%
\[
\mathcal{E}_{g}[I_{[W_{T}\geq-n]}+I_{[0\geq W_{T}\geq-n]}|\mathcal{F}%
_{t}]=\mathcal{E}_{g}[I_{[W_{T}\geq-n]}|\mathcal{F}_{t}]+\mathcal{E}%
_{g}[I_{[0\geq W_{T}\geq-n]}|\mathcal{F}_{t}].
\]

\end{lemma}

\begin{proof}
Let $W_{t,T}$ denote $W_{T}-W_{t}$. For each $a<b,$ it is easy to verify that
$I_{[W_{t,T}\geq a]}$ and $I_{[b\geq W_{t,T}\geq a]}$ are comonotonic. Then,
by $\mathcal{E}_{g}=\mathcal{C}_{g}$ and the comonotonic additivity of the
Choquet expectation, we have%
\begin{equation}\label{e2}
\mathcal{E}_{g}[I_{[W_{t,T}\geq a]}+I_{[b\geq W_{t,T}\geq a]}]=\mathcal{E}%
_{g}[I_{[W_{t,T}\geq a]}]+\mathcal{E}_{g}[I_{[b\geq W_{t,T}\geq a]}].
\end{equation}
On the other hand, for each $l_{1},l_{2}\in \mathbb{R},$ it is easy to show
that $f(x,y):=l_{1}I_{[x+y\geq-n]}+l_{2}I_{[0\geq x+y\geq-n]}$ satisfies the
condition in Remark \ref{re7}. Hence, we have%
\begin{equation}\label{e3}
\mathcal{E}_{g}[l_{1}I_{[W_{T}\geq-n]}+l_{2}I_{[0\geq W_{T}\geq-n]}%
|\mathcal{F}_{t}]=\mathcal{E}_{g}[l_{1}I_{[W_{t,T}\geq-n-\bar{a}]}%
+l_{2}I_{[-\bar{a}\geq W_{t,T}\geq-n-\bar{a}]}]|_{\bar{a}=W_{t}}.
\end{equation}
Combining (\ref{e3}) with (\ref{e2}) yields the result, and the
proof is complete.
\end{proof}

The following lemma is our main theorem with $d=1,$ which is an
extension of Lemma \ref{le3}.

\begin{lemma}\label{le9}
Suppose that $d=1$ and $g$ satisfies (H1)-(H3). Then $\mathcal{E}%
_{g}[\xi]=\mathcal{C}_{g}[\xi]$ for each $\xi \in L^{2}(\mathcal{F}_{T})$ if
and only if $g$ is independent of $y$ and is linear in $z$, i.e.,
$g(t,z)=g(t,1)z$ for almost every $t\in \lbrack0,T].$
\end{lemma}

\begin{proof}
If $g(t,z)=g(t,1)z,$ by the Girsanov Theorem, the $g$-expectation is a linear
mathematical expectation. Therefore, $\mathcal{E}_{g}[\xi]=\mathcal{C}_{g}%
[\xi]$ for each $\xi \in L^{2}(\mathcal{F}_{T})$ and the proof of
sufficient condition is complete. Now we prove the necessary
condition. By Lemma \ref{le5}, $g$ is independent of $y$. For each
$n\in \mathbb{N}$, consider the following
BSDEs:%
\begin{align*}
y_{t}^{n}  & =I_{[W_{T}\geq-n]}+I_{[0\geq W_{T}\geq-n]}+\int_{t}^{T}%
g(s,z_{s}^{n})ds-\int_{t}^{T}z_{s}^{n}dW_{s},\\
\tilde{y}_{t}^{n}  & =I_{[W_{T}\geq-n]}+\int_{t}^{T}g(s,\tilde{z}_{s}%
^{n})ds-\int_{t}^{T}\tilde{z}_{s}^{n}dW_{s},\\
\hat{y}_{t}^{n}  & =I_{[0\geq W_{T}\geq-n]}+\int_{t}^{T}g(s,\hat{z}_{s}%
^{n})ds-\int_{t}^{T}\hat{z}_{s}^{n}dW_{s}.
\end{align*}
By Lemma \ref{le8}, we have
$y_{t}^{n}=\tilde{y}_{t}^{n}+\hat{y}_{t}^{n}$ for each
$t\in \lbrack0,T].$ Form this, we have%
\begin{equation}\label{e4}
dP\times dt-a.s.,\quad g(t,\tilde{z}_{t}^{n}+\hat{z}_{t}^{n})=g(t,\tilde{z}%
_{t}^{n})+g(t,\hat{z}_{t}^{n}).
\end{equation}
On the other hand, it follows from Lemma \ref{le5} that $g$ is
positively homogeneous.
Thus we have for almost every $t\in \lbrack0,T],$%
\begin{equation}\label{e5}
g(t,z)=g(t,1)z^{+}+g(t,-1)z^{-},
\end{equation}
where $z^{+}=\max \{z,0\},$ $z^{-}=(-z)^{+}.$ Set $h(t):=g(t,1)+g(t,-1),$ by
(\ref{e4}) and (\ref{e5}), we have%
\begin{equation}\label{e6}
dP\times dt-a.s.,\quad
h(t)(\tilde{z}_{t}^{n}+\hat{z}_{t}^{n})^{-}=h(t)(\tilde
{z}_{t}^{n})^{-}+h(t)(\hat{z}_{t}^{n})^{-}.
\end{equation}
Also, $dP\times dt-a.s.$, $\tilde{z}_{t}^{n}=\frac{1}{\sqrt{2\pi(T-t)}}%
\exp(-\frac{(n+W_{t}+\int_{t}^{T}g(s,1)ds)^{2}}{2(T-t)})>0$ (see Lemma 8 in
\cite[Chen et al. (2005a)]{CCD}). This with (\ref{e6}) implies%
\begin{equation}\label{e7}
dP\times dt-a.s.,\quad h(t)I_{[\hat{z}_{t}^{n}<0]}=0.
\end{equation}
Let $(\bar{y}_{t},\bar{z}_{t})_{t\in \lbrack0,T]}$ denote the
solution of BSDE (\ref{e1}) corresponding to the generator $g$ and
terminal value $I_{[W_{T}\leq0]}.$ It follows from Lemma \ref{le1}
that $\hat{z}_{t}^{n}\rightarrow \bar{z}_{t}$ as $n\rightarrow
\infty$ in $L^{2}(0,T;\mathbb{R}).$ Thus we can choose
$n_{i}\rightarrow \infty$ such that $dP\times dt-a.s.$, $\hat{z}_{t}^{n_{i}%
}\rightarrow \bar{z}_{t}.$ Noting that $\bar{z}_{t}=\frac{-1}{\sqrt{2\pi(T-t)}%
}\exp(-\frac{(W_{t}-\int_{t}^{T}g(s,-1)ds)^{2}}{2(T-t)})<0,$ then by
(\ref{e7}), we can deduce that for almost every $t\in \lbrack0,T],$
$h(t)=0.$ Thus $g(t,z)=g(t,1)z$ for almost every $t\in \lbrack0,T].$
The proof is complete.
\end{proof}

\begin{corollary}\label{co10}
Suppose $g$ satisfies (H1)-(H3). If $\mathcal{E}_{g}[\xi]=\mathcal{C}_{g}%
[\xi]$ for each $\xi \in L^{2}(\mathcal{F}_{T})$, then $g$ is independent of
$y$ and is homogeneous in $z$, i.e., for almost every $t\in \lbrack0,T],$
$g(t,\lambda z)=\lambda g(t,z)$ for each $\lambda \in \mathbb{R}$.
\end{corollary}

\begin{proof}
For each fixed $a\in \mathbb{R}^{d}$ with $|a|=1$, set $\tilde{W}_{t}:=a\cdot
W_{t}$ and $\mathcal{\tilde{F}}_{t}:=\sigma \{ \tilde{W}_{s}:s\leq t\}$ for each
$t\in \lbrack0,T]$. Obviously, $(\tilde{W}_{t})_{t\in \lbrack0,T]}$ is a
$1$-dimensional Brownian motion. Define $\tilde{g}:[0,T]\times \mathbb{R\times
R\rightarrow R}$ by $\tilde{g}(t,y,z):=g(t,y,az)$. It is easy to verify that
$\tilde{g}$ satisfies (H1)-(H3). For each $\xi \in L^{2}(\mathcal{\tilde{F}%
}_{T})$, let $(y_{t},z_{t})_{t\in \lbrack0,T]}$ denote the solution of the
following BSDE:%
\[
y_{t}=\xi+\int_{t}^{T}\tilde{g}(s,y_{s},z_{s})ds-\int_{t}^{T}z_{s}d\tilde
{W}_{s},\ 0\leq t\leq T.
\]
Then it is easy to check that $(y_{t},az_{t})_{t\in \lbrack0,T]}$ is
the solution of BSDE (\ref{e1}) corresponding to the generator $g$
and terminal value
$\xi$. From this, we deduce that $\mathcal{E}_{g}[\xi]=\mathcal{E}_{\tilde{g}%
}[\xi]$ for each $\xi \in L^{2}(\mathcal{\tilde{F}}_{T})$. Noting that
$\mathcal{E}_{g}=\mathcal{C}_{g}$, we then get $\mathcal{E}_{\tilde{g}}%
[\xi]=\mathcal{C}_{\tilde{g}}[\xi]$ for each $\xi \in L^{2}(\mathcal{\tilde{F}%
}_{T})$. By Lemma \ref{le9}, we obtain
$\tilde{g}(t,y,z)=\tilde{g}(t,0,1)z$ for almost every $t\in
\lbrack0,T]$. Hence, by the Lipschitz assumption (H2), we have for
almost every $t\in \lbrack0,T]$, $g(t,y,\lambda a)=\lambda g(t,0,a)$
for each $\lambda \in \mathbb{R}$ and $a\in \mathbb{R}^{d}$ with
$|a|=1$, which implies that $g$ is independent of $y$ and is
homogeneous in $z$. We complete the proof.
\end{proof}

To prove the main theorem, we also pay more attention to the
following two lemmas.

\begin{lemma}\label{le11}
Suppose that $d=2$ and $g$ satisfies (H1)-(H3). If $\mathcal{E}_{g}%
[\xi]=\mathcal{C}_{g}[\xi]$ for each $\xi \in L^{2}(\mathcal{F}_{T})$, then for
each $\lambda \in \lbrack0,1]$, $t\in \lbrack0,T],$ $n\in \mathbb{N},$ we have%
\[
\mathcal{E}_{g}[I_{[W_{T}^{1}\geq n]}+\lambda I_{[W_{T}^{2}\geq0]}%
|\mathcal{F}_{t}]=\lambda \mathcal{E}_{g}[I_{[W_{T}^{1}\geq n]}+I_{[W_{T}%
^{2}\geq0]}|\mathcal{F}_{t}]+(1-\lambda)\mathcal{E}_{g}[I_{[W_{T}^{1}\geq
n]}|\mathcal{F}_{t}],
\]
where $W_{t}^{i}$ is the $i$-th component of $W_{t}$ with $i=1,2.$
\end{lemma}

\begin{proof}
Let $W_{t,T}^{i}$ denote $W_{T}^{i}-W_{t}^{i}$ with $i=1,2$. For
each fixed $\lambda \in \lbrack0,1],$ $a,b\in \mathbb{R}$, it is
easy to check that $(1-\lambda)I_{[W_{t,T}^{1}\geq a]}$ and
$\lambda(I_{[W_{t,T}^{1}\geq a]}+I_{[W_{t,T}^{2}\geq b]})$ are
comonotonic. The rest of the proof runs as in Lemma \ref{le8}, and
the proof is complete.
\end{proof}

\begin{lemma}\label{le12}
Suppose that $d=2$ and $g$ satisfies (H1)-(H3). If $\mathcal{E}_{g}%
[\xi]=\mathcal{C}_{g}[\xi]$ for each $\xi \in L^{2}(\mathcal{F}_{T})$, then $g$
is independent of $y$ and is linear in $z$, i.e., $g(t,z_{1},z_{2}%
)=g(t,1,0)z_{1}+g(t,0,1)z_{2}$ for almost every $t\in \lbrack0,T].$
\end{lemma}

\begin{proof}
It follows from Lemma \ref{le5} that $g$ is independent of $y$. For
each fixed
$\lambda \in(0,1),$ $n\in \mathbb{N},$ consider the following BSDEs:%
\begin{align*}
y_{t}^{\lambda,n} &  =I_{[W_{T}^{1}\geq n]}+\lambda I_{[W_{T}^{2}\geq0]}%
+\int_{t}^{T}g(s,z_{1,s}^{\lambda,n},z_{2,s}^{\lambda,n})ds-\int_{t}%
^{T}z_{1,s}^{\lambda,n}dW_{s}^{1}-\int_{t}^{T}z_{2,s}^{\lambda,n}dW_{s}^{2},\\
\tilde{y}_{t}^{n} &  =I_{[W_{T}^{1}\geq n]}+I_{[W_{T}^{2}\geq0]}+\int_{t}%
^{T}g(s,\tilde{z}_{1,s}^{n},\tilde{z}_{2,s}^{n})ds-\int_{t}^{T}\tilde{z}%
_{1,s}^{n}dW_{s}^{1}-\int_{t}^{T}\tilde{z}_{2,s}^{n}dW_{s}^{2},\\
\hat{y}_{t}^{n} &  =I_{[W_{T}^{1}\geq n]}+\int_{t}^{T}g(s,\hat{z}_{1,s}%
^{n},\hat{z}_{2,s}^{n})ds-\int_{t}^{T}\hat{z}_{1,s}^{n}dW_{s}^{1}-\int_{t}%
^{T}\hat{z}_{2,s}^{n}dW_{s}^{2}.
\end{align*}
By Lemma \ref{le11}, we have $y_{t}^{\lambda,n}=\lambda \tilde{y}_{t}^{n}%
+(1-\lambda)\hat{y}_{t}^{n}$ for each $t\in \lbrack0,T]$. From this, we deduce
that $dP\times dt-a.s.$,
\[
g(t,\lambda \tilde{z}_{1,t}^{n}+(1-\lambda)\hat{z}_{1,t}^{n},\lambda \tilde
{z}_{2,t}^{n}+(1-\lambda)\hat{z}_{2,t}^{n})=\lambda g(t,\tilde{z}_{1,t}%
^{n},\tilde{z}_{2,t}^{n})+(1-\lambda)g(t,\hat{z}_{1,t}^{n},\hat{z}_{2,t}^{n}).
\]
Since $\lambda \in(0,1)$ is arbitrary, by Lemma \ref{le5}, we obtain
that
$dP\times dt-a.s.$,%
\begin{equation}\label{e8}
g(t,\tilde{z}_{1,t}^{n}+l\hat{z}_{1,t}^{n},\tilde{z}_{2,t}^{n}+l\hat{z}%
_{2,t}^{n})=g(t,\tilde{z}_{1,t}^{n},\tilde{z}_{2,t}^{n})+g(t,l\hat{z}%
_{1,t}^{n},l\hat{z}_{2,t}^{n})\ \mbox{for each}\ l\geq0.
\end{equation}
Noting that $g(t,z_{1},0)=g(t,1,0)z_{1}$ for almost every $t\in \lbrack0,T],$
then we have%
\begin{equation}\label{e9}
dP\times dt-a.s.,\ (\hat{z}_{1,t}^{n},\hat{z}_{2,t}^{n})=(\frac{1}{\sqrt
{2\pi(T-t)}}\exp(-\frac{(n-W_{t}^{1}-\int_{t}^{T}g(s,1,0)ds)^{2}}{2(T-t)}),0).
\end{equation}
Combining (\ref{e8}) with (\ref{e9}), we get
\begin{equation}\label{e10}
dP\times dt-a.s.,\ g(t,\tilde{z}_{1,t}^{n}+p,\tilde{z}_{2,t}^{n}%
)=g(t,\tilde{z}_{1,t}^{n},\tilde{z}_{2,t}^{n})+g(t,p,0)\ \mbox{for
each}\ p\geq0.
\end{equation}
Let $(\bar{y}_{t},\bar{z}_{1,t},\bar{z}_{2,t})_{t\in \lbrack0,T]}$
be the solution of BSDE (\ref{e1}) corresponding to the generator
$g$ and terminal value $I_{[W_{T}^{2}\geq0]}$. By Lemma \ref{le1},
we have $(\tilde{z}_{1,t}^{n},\tilde
{z}_{2,t}^{n})\rightarrow(\bar{z}_{1,t},\bar{z}_{2,t})$ in $L^{2}%
(0,T;\mathbb{R}^{2}).$ Since $g$ satisfies Lipschitz assumption (H2), we get
for each $p\geq0,$%
\[
g(t,\tilde{z}_{1,t}^{n}+p,\tilde{z}_{2,t}^{n})\rightarrow g(t,\bar{z}%
_{1,t}+p,\bar{z}_{2,t})\ \mbox{in}\ L^{2}(0,T;\mathbb{R}).
\]
This with (\ref{e10}) implies that
\begin{equation}\label{e11}
dP\times dt-a.s.,\ g(t,\bar{z}_{1,t}+p,\bar{z}_{2,t})=g(t,\bar{z}_{1,t}%
,\bar{z}_{2,t})+g(t,p,0)\ \mbox{for each}\ p\geq0.
\end{equation}
Also, we have%
\begin{equation}\label{e12}
dP\times dt-a.s.,\ (\bar{z}_{1,t},\bar{z}_{2,t})=(0,\frac{1}{\sqrt{2\pi(T-t)}%
}\exp(-\frac{(W_{t}^{2}+\int_{t}^{T}g(s,0,1)ds)^{2}}{2(T-t)})).
\end{equation}
It follows from (\ref{e11}), (\ref{e12}) and Lemma \ref{le5} that
for almost every $t\in
\lbrack0,T],$%
\begin{equation}\label{e13}
g(t,p,1)=g(t,0,1)+g(t,p,0)\ \mbox{for each}\ p\geq0.
\end{equation}
From (\ref{e13}) and Corollary \ref{co10}, we can easily deduce that
for almost every
$t\in \lbrack0,T],$%
\begin{equation}\label{e14}
g(t,z_{1},z_{2})=g(t,1,0)z_{1}+g(t,0,1)z_{2}\ \mbox{for
each}\ z_{1}\cdot z_{2}%
\geq0.
\end{equation}
On the other hand, set $\tilde{W}_{t}:=(W_{t}^{1},-W_{t}^{2})$ and $\tilde
{g}(t,z_{1},z_{2})=g(t,z_{1},-z_{2}).$ Analysis similar to that in the proof
of Corollary \ref{co10} shows that $\mathcal{E}_{\tilde{g}}[\xi]=\mathcal{C}%
_{\tilde{g}}[\xi]$ for each $\xi \in L^{2}(\mathcal{F}_{T})$. Then
we have (\ref{e14})
for $\tilde{g}$, which gives that for almost every $t\in \lbrack0,T],$%
\[
g(t,z_{1},z_{2})=g(t,1,0)z_{1}+g(t,0,1)z_{2}\ \mbox{for
each}\ z_{1}\cdot z_{2}%
\leq0.
\]
The proof is now complete.
\end{proof}

We now prove the main theorem.

\textbf{Proof of Theorem \ref{th4}. }The sufficient condition can be
proved by the same method as in Lemma \ref{le9}. We only prove the
necessary condition. For $d=2,$ by Lemma \ref{le12}, the result
holds. We only prove the case $d>2.$ For each fixed
$a\in \mathbb{R}^{d-1}$ with $|a|=1,$ set $\tilde{W}_{t}:=(a\cdot(W_{t}%
^{1},\ldots,W_{t}^{d-1}),W_{t}^{d})$ and $\mathcal{\tilde{F}}_{t}%
:=\sigma \{ \tilde{W}_{s}:s\leq t\}$ for each $t\in \lbrack0,T]$. By
Lemma \ref{le5}, $g$ is independent of $y$, we define
\thinspace$\tilde{g}:[0,T]\times
\mathbb{R\times R\rightarrow R}$ by $\tilde{g}(t,z_{1},z_{2}):=g(t,az_{1}%
,z_{2})$. As in the proof of Corollary \ref{co10}, we can get $\mathcal{E}_{\tilde{g}%
}[\xi]=\mathcal{C}_{\tilde{g}}[\xi]$ for each $\xi \in
L^{2}(\mathcal{\tilde
{F}}_{T})$. By Lemma \ref{le12}, we have for almost every $t\in \lbrack0,T],$%
\[
\tilde{g}(t,z_{1},z_{2})=\tilde{g}(t,1,0)z_{1}+\tilde{g}(t,0,1)z_{2}.
\]
Since $a$ is arbitrary, by Corollary \ref{co10}, we obtain for
almost every
$t\in \lbrack0,T],$%
\[
g(t,z_{1},\ldots,z_{d-1},z_{d})=g(t,z_{1},\ldots,z_{d-1},0)+g(t,e_{d})z_{d}.
\]
Define $\bar{g}:[0,T]\times \mathbb{R}^{d-1}\mathbb{\rightarrow R}$ by $\bar
{g}(t,z):=g(t,z,0).$ We now apply the above argument again, with $g$ replaced
by $\bar{g}$, to obtain that for almost every $t\in \lbrack0,T],$%
\[
\bar{g}(t,z_{1},\ldots,z_{d-2},z_{d-1})=\bar{g}(t,z_{1},\ldots,z_{d-2}%
,0)+\bar{g}(t,0,\ldots,0,1)z_{d-1},
\]
that is%
\[
g(t,z_{1},\ldots,z_{d-2},z_{d-1},0)=g(t,z_{1},\ldots,z_{d-2},0,0)+g(t,e_{d-1}%
)z_{d-1}.
\]
Continuing this process, we can prove that $g(t,z)=\sum_{i=1}^{d}%
g(t,e_{i})z_{i}$ for almost every $t\in \lbrack0,T].$ The proof is complete.
$\Box$

\end{document}